what is the square root of 3^2 + 10times(the square root 16) p.s. It is one big square root, and 16 is a square root inside of the big square root
Found 3 solutions by Theo, MathTherapy, ikleyn:
Answer by Theo(13342)

you are looking at sqrt(3^2 + 10 * sqrt(16))
within your square root sign, 3^2 = 9 and 10 * sqrt(16) = 10 * 4 = 40
that makes it sqrt(9 + 40) which is sqrt(49) which is equal to 7.

Answer by MathTherapy(10551)

 what is the square root of 3^2 + 10times(the square root 16) p.s. It is one big square root, and 16 is a square root inside of the big square root


Answer by ikleyn(52777)

.

By posting these simple assignments to the forum instead of doing them on yourself,
you simply do everything against your own interests and against your own development.